OpenStudy (anonymous):

Which shows 4x + 2y = -16 in slope-intercept form? 1. y = 2x -8 2. y = -2x -8 3. y = 1/2x -8 4. y = 1/2x -16 Which pair of lines is perpendicular? 1. 10y = -9 and y = -9/10x +6 2. -4x +5y = -5 and y = -4/5x +7 3. -8/3x +9 = y and 8y = 3x -32 4. y = -5x +2 and y + -5x -1/2 A line has a slope of -3 and passes through the point (5, -1). What is the equation of the line? 1. -3x + y= -16 2. -3x + y= -2 3. 3x + y= 2 4. 3x + y= 14

OpenStudy (dayakar):

do u have any idea

OpenStudy (anonymous):

no @dayakar

OpenStudy (dayakar):

1)slope intercept form is y=mx+b [here m is slope and b is y-intercept] we have to change the equation into above form can u do it

OpenStudy (dayakar):

step 1)divide each term with 2 step 2) add -2x both sides u can get the answer

OpenStudy (dayakar):

2) if product of slopes of two lines = -1 then the two lines are perpendicular

OpenStudy (dayakar):

3)equation of the line whose slope (m) = -3 and passing through the point (x1 ,y1) = (5, -1) is y- y1 = m (x- x1) replace m, x1, y1 with given values and simplify it